A ball is projected so as to pass a wall at a distance 'a' from the point of projection at an angle of 45 degrees and falls at a distance 'b' on the other side of the wall.


If 'h' is the height of the wall....


h =


A) a(2^1/2)


B) b(2^1/2)


C) [(2^1/2)ab]/(a+b)


D) (ab)/(a+b)

Grade:12

1 Answers

Debraj Basu
12 years ago

Rquation of trajectory, y=xtan45(1-(x/R))

h=a(1-(a/a+b))=ab/(a+b)